OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) =(x+9)/(x^2+8x+5)

OpenStudy (anonymous):

the degree of the numerator is 1, the degree of the denominator is 2 since the denominator has a larger degree, the horizontal asymptote is \(y=0\)

OpenStudy (anonymous):

to find the vertical asymptotes, set \(x^2+8x+5=0\) and solve for \(x\)

OpenStudy (anonymous):

this one doesn't factor, but you can complete the square easily \[x^2+8x+5=0\] \[x^2+8x=-5\] \[(x+4)^2=-5+16=11\] \[x+4=\pm\sqrt{11}\] \[x=-4\pm\sqrt{11}\]

OpenStudy (anonymous):

look at first post, answer is there